OpenStudy (anonymous):

13+3g-2(-5+g) I need help with how to solve this. Please show all steps!

OpenStudy (anonymous):

13 + 3g - 2(-5 + g) = 13 + 3g +10 - 2g = = 23 - g

OpenStudy (anonymous):

First of all do whats in the ( )

OpenStudy (anonymous):

first of all Welcome To Open Study

OpenStudy (anonymous):

Thank you Aiko! So, first I'd multiply -2 to both things in parenthesis?

OpenStudy (dan815):

yes

OpenStudy (dan815):

did you learn PEDMAS

OpenStudy (anonymous):

Yes I did.

OpenStudy (dan815):

okay yes you just follow that order

OpenStudy (dan815):

if you see that the parenthesis cannot be simplified more*, you move onto the next one in PEDMAS

OpenStudy (anonymous):

On the second line wouldn't it be +10?

OpenStudy (dan815):

very good :)

OpenStudy (anonymous):

Thanks. :) Would I add 2g to both sides now?

OpenStudy (dan815):

lol what you are done

OpenStudy (dan815):